OpenStudy (anonymous):

Mount Everest is approx. 9km high. Calculate how much less a mountaineer of mass 100kg(with a back pack would weigh at its summit, compares to her weight at sea level.

OpenStudy (anonymous):

we will first find the gravitational force by F=GM/r^2, the use that to find the weight, w=mg right?

OpenStudy (jamesj):

Use Newton's Universal Law of Gravity as you did in your last question. You'll need M = mass of earth R = radius of earth at sea level m = her weight = 100 kg Then her weight at sea level is \[Weight_{sea \ level} = F_{gravity \ @ \ sea \ level} = \frac{GMm}{R^2} \] while atop Everest her weight is \[ Weight_{Everest} = F_{gravity \ atop \ Everest} = \frac{GMm}{(R + 9)^2} \]

OpenStudy (jamesj):

actually, you should measure distance in meters, so you need R in meters, and then atop Everest, the force due to gravity is \[ \frac{GMm}{(R+9000)^2} \]

OpenStudy (anonymous):

the equation I wrote should be g=GM/r2 Isn't it more simpler to use this equation than GMm/r2?

OpenStudy (jamesj):

As weight, W = mg, you're doing exactly the same thing. Just breaking up the calculation into two pieces.

OpenStudy (anonymous):

true :)